# U-MARVEL: Unveiling Key Factors for Universal Multimodal Retrieval via Embedding Learning with MLLMs

### v1

@@ 5 Conclusion and Limitations
Despite the promising results achieved by U-MARVEL, there remain several limitations that can be further improved. First, it primarily focuses on text and image modalities, leaving the inclusion of other modalities (e.g., audio) for future work. Second, while effective in zero-shot scenarios, its integration with LLMs in retrieval-augmented generation (RAG) applications remains underexplored. Lastly, experiments were limited to 7B-sized models due to submission time constraints; future work will extend support models of various sizes, enabling flexible trade-offs between performance and cost.
@@ Appendix A M-BEIR Dataset & Metrics
Datasets:
We evaluate the retrieval performance of our method with the M-BEIR [51] benchmark. This dataset consist of 8 possible multimodal retrieval tasks involving texts and images. It is constructed based on 10 different datasets from 4 domains. Moreover, according to the datasets and task types (i.e., the modalities of queries and candidates), the multi-modal retrieval tasks are further subdivided into 16 types.
@@
| VisualNews [33] | Identify news-related image match with the description | News | 99K | 20K | 20K | 542K | |
| MSCOCO [31] | Find an everyday image match with caption | Misc. | 100K | 24.8K | 24.8K | 5K | |
| Fashion200K [16] | Based on fashion description, retrieve matched image | Fashion | 15K | 1.7K | 1.7K | 201K | |
| WebQA [5] | Find an paragraph from Wikipedia to answer the question | Wiki | 16K | 1.7K | 2.4K | 544K | |
| EDIS [35] | Find a news image matching with the caption | News | 26K | 3.2K | 3.2K | 1M | |
| WebQA [5] | Find a Wiki image that answer the question | Wiki | 17K | 1.7K | 2.5K | 403K | |
| VisualNews [33] | Provide a news-related caption for the displayed image | News | 100K | 20K | 20K | 537K | |
| MSCOCO [31] | Find a caption describe the an image | Misc. | 113K | 5K | 5K | 25K | |
| Fashion200K [16] | Find a description for the fashion item in the image | Fashion | 15K | 4.8K | 4.8K | 61K | |
| NIGHTS [12] | Find an image that is identical to the given image | Misc. | 16K | 2K | 2K | 40K | |
| OVEN [20] | Retrieve a Wiki text that answer the given query about the image | Wiki | 150K | 50K | 50K | 676K | |
| InfoSeek [7] | Find an article that answers the given question about the image | Wiki | 141K | 11K | 11K | 611K | |
| FashionIQ [52] | Find an image to match the fashion image and style note | Fashion | 16K | 2K | 6K | 74K | |
| CIRR [38] | I’m looking for a similar everyday image with the described changes | Misc. | 26K | 2K | 4K | 21K | |
| OVEN [20] | Find a Wiki image-text pair to answer a question regarding an image | Wiki | 157K | 14.7K | 14.7K | 335K | |
| InfoSeek [7] | Find a Wiki image-text pair to answers my question about this image | Wiki | 143K | 17.6K | 17.6K | 481K | |
@@
Metrics:
Following the evaluation protocols by M-BEIR benchmark, we use the Recall@10 (R@10) for Fashion200K [16] and FashionIQ [52], and the Recall@5 (R@5) for the remaining datasets. We report the average score of all test queries as the criterion for measuring the retrieval accuracy.
@@ Appendix B Zero-shot Datasets & Metrics
Datasets:
We have examined the zero-shot capabilities of U-MARVEL in retrieval tasks, specifically focusing on image-text cross-modal retrieval and text-video retrieval. For tasks of cross-modal retrieval between images and texts, we use the following datasets: ShareGPT4V [21], Urban-1K [59], CIRCO [3], Flickr [40], GeneCIS [46], Visual Dialog [11], Multi-round FashionIQ [57], CC-Neg [44], and Sugar-Crepe [18]. For text-video retrieval, the datasets used are MSR-VTT [53] and MSVD [6]. We present the details of the Unseen Dataset in Table 11. Although some of them are actually adapted from MSCOCO or FashionIQ, we still treat these datasets as unseen datasets due to the significant differences in their captions or query formats, following [36]. For instance, the captions in Urban1K consist of extended captions generated by GPT-4V [21], while the query format of CIRCO combines a reference image with a relative caption. These differences create a substantial disparity compared to the original COCO [31] dataset. Table 11 provides a detailed presentation of the information about the zero-shot datasets.
@@
The first row of Table 8 indicates the retrieval task type: for text queries, for image queries, for dialog queries, for multi-interleaved image-text queries, for text candidates, for image candidates, and ITM for the Image-Text Matching task.
@@
| ShareGPT4V [21] | 1K | 1K | |
| Urban-1K [59] | 1K | 1K | |
| Flickr30K [40] | 1K | 5K | |
| CIRCO [3] | 800 | 120K | |
| GeneCIS [46] | 8K | 10 - 15 | |
| Visual Dialog [11] | 2K | 2K | |
| Multi-round FashionIQ [57] | 2.4K | 6.2K | |
| CC-Neg [44] | 40K | 2 | |
| Sugar-Crepe [18] | 7.5K | 2 | |
| MSR-VTT [53] | 670 | 27K | |
| MSVD [6] | 1K | 1K |
@@ Appendix C Implementation Details
We select Qwen2-VL-7B-Instruct [47] as the pretrained MLLM backbone to instantiate both of the retriever and the reranker. Throughout all experiments, we froze the visual side and employed LoRA [19] to fine-tune the LLM component of the model using Ascend-910B GPUs. For embedding models, the InfoNCE loss [39] is selected as our training objective:
where , , represent the embeddings of the query, the positive candidate, and any sample from the candicate set, respectively. is the temperature parameter. means the cosine similarity.
For reranking models, we train point-wise rerankers to classify whether a pair of query and candidate items is relevant. The loss function for the reranking model is formally defined as follows:
where if the candidate belongs to the set of positive candidates, and otherwise. Detailed training configuration are shown in Table 12.
@@
where is a tunable weighting hyperparameter. Based on this final score, the candidate set is reranked to generate the final ranked list. We refer to this unified framework as . The detailed training configurations are provided in Table 12.
@@ Appendix D Supplemental Experimental Results
Table 14 presents a comprehensive comparison of our proposed U-MARVEL framework with state-of-the-art methods on the M-BEIR benchmark under the global pooling setting. The results demonstrate that, within the single-model category, U-MARVEL significantly outperforms leading approaches such as MM-Embed [30] and LamRA-Ret [36], by a large margin. Furthermore, when integrated into a recall-then-rerank pipeline, the enhanced variant achieves the best overall performance, further solidifying the effectiveness of our framework. These findings highlight the versatility and superiority of U-MARVEL across different retrieval paradigms, showcasing its potential for advancing state-of-the-art retrieval systems.
@@
To evaluate the cross-task generalization ability of the proposed method, we trained the model on the five tasks, reserving the remaining three tasks as held-out tasks for evaluation. As presented in Table D, our method achieves an average improvement of 2.4%/1.6% across the three held-out tasks compared with SoTA approaches.
@@
| Task | Dataset | Metric | Single Model | +Rerank | |||||
[51] |
[51] |
MM-Embed [30] |
LamRA-Ret [36] |
U-MARVEL |
LamRA [36] |
| VisualNews | R@5 | 23 | 42.6 | 41 | 41.3 | 47.2 | 46.9 | 48.8 | |
| MSCOCO | R@5 | 75.6 | 77.9 | 71.3 | 75.4 | 72.8 | 78 | 70.1 | |
| Fashion200K | R@10 | 25.4 | 17.8 | 17.1 | 28.7 | 33.3 | 32.5 | 33.8 | |
| WebQA | R@5 | 79.5 | 84.7 | 95.9 | 85.8 | 96.7 | 96.5 | 98.3 | |
| EDIS | R@5 | 50.3 | 59.4 | 68.8 | 62.5 | 78.7 | 74.4 | 80.8 | |
| WebQA | R@5 | 79.7 | 78.8 | 85 | 81 | 87.7 | 87.1 | 88.3 | |
| VisualNews | R@5 | 21.1 | 42.8 | 41.3 | 39.3 | 47.2 | 47.6 | 49.8 | |
| MSCOCO | R@5 | 88.8 | 92.3 | 90.1 | 90.4 | 93.5 | 92.4 | 86.0 | |
| Fashion200K | R@10 | 27.6 | 17.9 | 18.4 | 30.4 | 34.9 | 36.6 | 36.8 | |
| NIGHTS | R@5 | 33 | 32 | 32.4 | 32.1 | 34.0 | 34.2 | 34.8 | |
| OVEN | R@5 | 38.7 | 39.2 | 42.1 | 48.4 | 58.3 | 54 | 58.7 | |
| InfoSeek | R@5 | 19.7 | 24 | 42.3 | 48.7 | 52.2 | 58.7 | 56.9 | |
| FashionIQ | R@10 | 28.5 | 24.3 | 25.7 | 33.1 | 36.0 | 37.4 | 37.4 | |
| CIRR | R@5 | 51.4 | 43.9 | 50 | 50.5 | 56.0 | 59.7 | 58.4 | |
| OVEN | R@5 | 57.8 | 60.2 | 64.1 | 70 | 73.1 | 72.6 | 74.9 | |
| InfoSeek | R@5 | 27.7 | 44.6 | 57.7 | 60 | 69.2 | 74 | 73.8 | |
| Avg. | 45.5 | 48.9 | 52.7 | 54.9 | 60.7 | 61.4 | 61.8 | ||
@@
| UniIR-BLIP [51] | 33.0 | 41.0 | 22.4 | 55.8 | 33.0 | 37.0 |
| UniIR-CLIP [51] | 32.0 | 45.5 | 27.9 | 67.6 | 48.9 | 44.4 |

### v2

@@ 5 Conclusion and Limitations
Despite the promising results achieved by U-MARVEL, there remain several limitations that can be improved. First, it focuses on text and image modalities, leaving the inclusion of other modalities (e.g., audio) for future work. Second, while effective in zero-shot scenarios, its integration with LLMs in retrieval-augmented generation (RAG) applications remains underexplored. Lastly, experiments were limited to 7B-sized models due to submission time constraints; future work will extend support models of various sizes, enabling flexible trade-offs between performance and cost.
Ethics statement
This study strictly adheres to the full requirements of the Code of Ethics outlined by the International Conference on Learning Representations (ICLR) . We have thoroughly read and adhered to this code during the research design, data handling, experimental execution, and manuscript writing phases, ensuring that the entire research process aligns with academic integrity and ethical guidelines.
Reproducibility statement
To ensure the reproducibility of this study, we fixed the global random seed to 42 across Python, NumPy, and PyTorch environments. This configuration minimizes stochasticity in model initialization, data sampling, and training dynamics. Consequently, the results reported in Tables 1–9 and Tables 13–18 reflect deterministic runs, aligning with standard practices in large-scale model research where computational constraints render multi-seed training infeasible. To facilitate direct replication, the source code and datasets are publicly available via the GitHub repository linked in the Abstract. Additionally, detailed experimental environment configurations are included in the Appendix.
@@ Appendix A M-BEIR Dataset & Metrics
Datasets.
We evaluate the retrieval performance of our method with the M-BEIR (Wei et al., 2024) benchmark. This dataset consist of 8 possible multimodal retrieval tasks involving texts and images. It is constructed based on 10 different datasets from 4 domains. Moreover, according to the datasets and task types (i.e., the modalities of queries and candidates), the multi-modal retrieval tasks are further subdivided into 16 types.
@@
| VisualNews (Liu et al., 2020) | Identify news-related image match with the description | News | 99K | 20K | 20K | 542K | |
| MSCOCO (Lin et al., 2014) | Find an everyday image match with caption | Misc. | 100K | 24.8K | 24.8K | 5K | |
| Fashion200K (Han et al., 2017) | Based on fashion description, retrieve matched image | Fashion | 15K | 1.7K | 1.7K | 201K | |
| WebQA (Chang et al., 2022) | Find an paragraph from Wikipedia to answer the question | Wiki | 16K | 1.7K | 2.4K | 544K | |
| EDIS (Liu et al., 2023b) | Find a news image matching with the caption | News | 26K | 3.2K | 3.2K | 1M | |
| WebQA (Chang et al., 2022) | Find a Wiki image that answer the question | Wiki | 17K | 1.7K | 2.5K | 403K | |
| VisualNews (Liu et al., 2020) | Provide a news-related caption for the displayed image | News | 100K | 20K | 20K | 537K | |
| MSCOCO (Lin et al., 2014) | Find a caption describe the an image | Misc. | 113K | 5K | 5K | 25K | |
| Fashion200K (Han et al., 2017) | Find a description for the fashion item in the image | Fashion | 15K | 4.8K | 4.8K | 61K | |
| NIGHTS (Fu et al., 2023) | Find an image that is identical to the given image | Misc. | 16K | 2K | 2K | 40K | |
| OVEN (Hu et al., 2023) | Retrieve a Wiki text that answer the given query about the image | Wiki | 150K | 50K | 50K | 676K | |
| InfoSeek (Chen et al., 2023) | Find an article that answers the given question about the image | Wiki | 141K | 11K | 11K | 611K | |
| FashionIQ (Wu et al., 2021) | Find an image to match the fashion image and style note | Fashion | 16K | 2K | 6K | 74K | |
| CIRR (Liu et al., 2021) | I’m looking for a similar everyday image with the described changes | Misc. | 26K | 2K | 4K | 21K | |
| OVEN (Hu et al., 2023) | Find a Wiki image-text pair to answer a question regarding an image | Wiki | 157K | 14.7K | 14.7K | 335K | |
| InfoSeek (Chen et al., 2023) | Find a Wiki image-text pair to answers my question about this image | Wiki | 143K | 17.6K | 17.6K | 481K | |
@@
Metrics.
Following the evaluation protocols by M-BEIR benchmark, we use the Recall@10 (R@10) for Fashion200K (Han et al., 2017) and FashionIQ (Wu et al., 2021), and the Recall@5 (R@5) for the remaining datasets. We report the average score of all test queries as the criterion for measuring the retrieval accuracy.
@@ Appendix B Zero-shot Datasets & Metrics
Datasets.
We have examined the zero-shot capabilities of U-MARVEL in retrieval tasks, specifically focusing on image-text cross-modal retrieval and text-video retrieval. For tasks of cross-modal retrieval between images and texts, we use the following datasets: ShareGPT4V (Hurst et al., 2024), Urban-1K (Zhang et al., 2024a), CIRCO (Baldrati et al., 2023), Flickr (Plummer et al., 2015), GeneCIS (Vaze et al., 2023), Visual Dialog (Das et al., 2017), Multi-round FashionIQ (Yuan and Lam, 2021), CC-Neg (Singh et al., 2024), and Sugar-Crepe (Hsieh et al., 2023). For text-video retrieval, the datasets used are MSR-VTT (Xu et al., 2016) and MSVD (Chen and Dolan, 2011). We present the details of the Unseen Dataset in Table 11. Although some of them are actually adapted from MSCOCO or FashionIQ, we still treat these datasets as unseen datasets due to the significant differences in their captions or query formats, following (Liu et al., 2024b). For instance, the captions in Urban1K consist of extended captions generated by GPT-4V (Hurst et al., 2024), while the query format of CIRCO combines a reference image with a relative caption. These differences create a substantial disparity compared to the original COCO (Lin et al., 2014) dataset. Table 11 provides a detailed presentation of the information about the zero-shot datasets.
@@
The first row of Table 8 indicates the retrieval task type: for text queries, for image queries, for dialog queries, for multi-interleaved image-text queries, for text candidates, for image candidates, and ITM for the Image-Text Matching task. Datasets used in our experiments include ShareGPT4V (ST4V), Urban-1K* (U-1K*), Flickr, CIRCO* (CIRCO*), GeneCIS* (GCIS*), Visual Dialog* (V-Dia*), Multi-round FashionIQ* (M-FIQ*), CC-Neg (C-Neg), and Sugar-Crepe* (S-Cre*) in Table 8. * indicates that the images in these datasets are sourced from COCO or FashionIQ.
@@
| ShareGPT4V (Hurst et al., 2024) | 1K | 1K | |
| Urban-1K (Zhang et al., 2024a) | 1K | 1K | |
| Flickr30K (Plummer et al., 2015) | 1K | 5K | |
| CIRCO (Baldrati et al., 2023) | 800 | 120K | |
| GeneCIS (Vaze et al., 2023) | 8K | 10 - 15 | |
| Visual Dialog (Das et al., 2017) | 2K | 2K | |
| Multi-round FashionIQ (Yuan and Lam, 2021) | 2.4K | 6.2K | |
| CC-Neg (Singh et al., 2024) | 40K | 2 | |
| Sugar-Crepe (Hsieh et al., 2023) | 7.5K | 2 | |
| MSR-VTT (Xu et al., 2016) | 670 | 27K | |
| MSVD (Chen and Dolan, 2011) | 1K | 1K |
@@ Appendix C Implementation Details
We select Qwen2-VL-7B-Instruct (Wang et al., 2024a) as the pretrained MLLM backbone to instantiate both of the retriever and the reranker. Throughout all experiments, we froze the visual side and employed LoRA (Hu et al., 2022) to fine-tune the LLM component of the model. For embedding models, the InfoNCE loss (Oord et al., 2018) is selected as our training objective:
where , , represent the embeddings of the query, the positive candidate, and any sample from the candidate set, respectively. is the temperature parameter. means the cosine similarity.
Recall-Rerank Model.
Some studies (Liu et al., 2024b; Lin et al., 2024) further refine retrieval results by applying a reranking step on top of the recall model. Inspired by (Liu et al., 2024b), we randomly sample top- hard negatives from the training hard negative stage to construct data triplets (query, positive, hard negative). Based on the progressive transition model, we perform supervised fine-tuning of the model shown in Figure 1(b) using the loss function defined in Eq. 4. Detailed training configuration are shown in Table 12.
where if the candidate belongs to the set of positive candidates, and otherwise.
During inference, we adopt the reranker’s predicted probability of the token “YES” as the relevance score, as shown in Eq. 5. After training the reranker, we employ the hard negative mining model for recall, obtaining the top- candidates along with their corresponding retrieval scores for each query. These candidates are then evaluated using the reranker model, which produces scores . The final ranking is generated by combining both sets of scores through Eq. 6. This completes our recall-then-rerank pipeline, producing optimized query-candidate matching scores.
where is set to 0.5.
Improved Distillation.
For the distillation stage, we use the following loss function:
where represents the multi-model ensemble scores and the single model’s predictions. The scores are softmax-normalized before computing the KL divergence.
@@
where is a weighting hyperparameter, set to 0.5. Based on this final score, the candidate set is reranked to generate the final ranked list. We refer to this unified framework as . The detailed training configurations are provided in Table 12.
@@ Appendix D Supplemental Experimental Results
Table 17 presents a comprehensive comparison of our proposed U-MARVEL framework with state-of-the-art methods on the M-BEIR benchmark under the global pooling setting. The results demonstrate that, within the single-model category, U-MARVEL significantly outperforms leading approaches such as MM-Embed (Lin et al., 2024) and LamRA-Ret (Liu et al., 2024b), by a large margin. Furthermore, when integrated into a recall-then-rerank pipeline, the enhanced variant achieves the best overall performance, further solidifying the effectiveness of our framework. These findings highlight the versatility and superiority of U-MARVEL across different retrieval paradigms, showcasing its potential for advancing state-of-the-art retrieval systems.
@@
To evaluate the cross-task generalization ability of the proposed method, we trained the model on the five tasks, reserving the remaining three tasks as held-out tasks for evaluation. As presented in Table 18, our method achieves an average improvement of 2.4%/1.6% across the three held-out tasks compared with SoTA approaches.
Analysis on Training Data for Progressive Transition.
Our method adopts a progressive training pipeline: NLI CC3M M-BEIR, achieving scores of 57.7%/55.8% as shown in Table 3. During the progressive training phase, we attempted to incorporate high-quality image-text pairs from ShareGPT4V, a dataset characterized by highly detailed image captions. However, contrary to expectations, utilizing ShareGPT4V data—either independently or mixed with CC3M—resulted in performance degradation on the M-BEIR benchmark, as demonstrated in Table 13. We hypothesize that this decline stems from a significant distribution shift: the dense, exhaustive descriptive text in ShareGPT4V contrasts sharply with the concise, query-oriented text typical of retrieval tasks (e.g., M-BEIR). This discrepancy likely hinders the model’s alignment with the target retrieval objective. Furthermore, we explored an alternative progressive path: NLI TART (Asai et al., 2023) M-BEIR, which also led to a substantial decrease in performance. We conjecture that the diverse text instruction tasks within TART interfere with the model’s effective alignment for multimodal retrieval.
Analysis on Learnable Temperature of InfoNCE.
We analyze the impact of the learnable temperature parameter , initialized at 0.05 in Progressive Transition (Stage-3).
Superior Performance over Fixed : Learnable consistently outperforms static settings. In Table 4, ID-7 (batch size 3840) and ID-8 (batch size 7680) achieve scores of 60.1% and 60.3%, respectively, surpassing the fixed baseline (ID-4 and ID-6, both at 58.9%). We further conducted an experiment fixing (aligned with the converged value of ID-7), which yielded a suboptimal score of 59.8%. This comparison confirms that dynamic adjustment adapts better to the optimization landscape than even a carefully manually tuned static value.
Adaptive Convergence: We observe stable convergence of across different batch sizes. Specifically, we observed convergence to 0.0206 for batch size 3840 and 0.0195 for batch size 7680. The slight decrease in the larger batch setting aligns with findings in (Sun and Li, 2024), where a sharper distribution is necessitated to handle increased batch noise.
Stage-wise Evolution: As detailed in Table 12, fluctuates adaptively across stages (Stage-3: 0.0206 Hard-neg: 0.0188 Distillation: 0.0204). Notably, drops during Hard-negative Mining to increase sensitivity to hard gradients.
Analysis on Generalization Ability of U-MARVEL.
To investigate the generalization ability of U-MARVEL across different MLLM backbones, we extended our framework to the Qwen3-VL-4B model. As presented in Table 16, the experimental results align consistently with the conclusions drawn from the 7B model, showing steady performance improvements across the three training stages—Progressive Transition, Hard Negative Mining, and Distillation. This validates that our proposed training recipe is robust and agnostic to specific model architectures. Remarkably, despite having significantly fewer parameters, the Qwen3-VL-4B model trained with U-MARVEL achieves a Local Average Recall of 58.8% and a Global Average Recall of 56.2% on the M-BEIR benchmark. These results not only demonstrate the effectiveness of our method but also establish a new state-of-the-art, surpassing the previous best-performing method, LamRA-Ret (Liu et al., 2024b) (Local: 56.6%, Global: 54.9%) with a larger 7B backbone. This underscores the strong generalization potential of U-MARVEL in enabling smaller models to achieve superior retrieval capabilities in both local and global search scenarios.
| ID | Training Pipeline | Local Avg. | Global Avg. |
| 0 | NLI ShareGPT4V M-BEIR | 56.9 | 54.0 |
| 1 | NLI CC3M&ShareGPT4V M-BEIR | 57.2 | 55.2 |
| 2 | NLI TART M-BEIR | 50.3 | 36.6 |
Appendix E Distillation Computational Cost Analysis
The distillation process consists of two main parts: (1) The teacher model infers a similarity matrix; (2) The student model is trained to extract features, compute scores, and learn from the teacher model’s similarity matrix. In terms of computational cost, the recall stage primarily focuses on extracting the feature representations of the samples. On the other hand, the rerank stage is primarily concerned with calculating the similarity scores between the query and candidates. Despite the differences in tasks, the inference time for a single pass through both stages is approximately the same. Moreover, we assume that during the distillation phase, the training time for a feature by the student model is approximately twice that of inference, as it includes both forward and backward propagation. Table 14 presents a detailed comparison between the traditional and improved distillation methods. We adopt the following notation for our analysis:
: The total number of queries in the dataset (assuming a ratio of 1:1 with positive samples, the candidate pool size is also ).
: The number of queries in a single batch.
: The number of top- hard negatives selected in our improved method.
E.1 Teacher Model Inference Stage
This stage is subdivided into Recall and Rerank phases.
Recall:
Both the traditional and improved methods require feature extraction for all queries and candidate documents to compute retrieval scores. With queries and positive candidates, the total inference cost is . For the improved method, the additional overhead of nearest neighbor search to identify top- hard negatives is computationally negligible compared to feature extraction and is therefore omitted.
Rerank:
This phase focuses on computing query-candidate similarity scores.
Traditional Method: Computes scores for all pairs within a batch (utilizing in-batch negatives), resulting in an similarity matrix per batch. The total cost is .
Improved Method: Computes scores strictly for the query, the positive sample, and the mined hard negatives, resulting in an matrix per batch. The total cost is .
E.2 Student Model Training Stage
Traditional Method: The student model processes the query and positive sample pairs, totaling features. Applying the training cost multiplier (factor of 2 for forward/backward passes), the total training cost is .
Improved Method: The input augments the query and positive sample with hard negatives, totaling features. Consequently, the training cost is .
E.3 Overall Complexity
Aggregating the costs from both stages, the total computational complexity is for the traditional method and for the improved method. Given that the batch size is typically much larger than the number of hard negatives (i.e., ), our improved method significantly reduces the computational burden, particularly during the Rerank phase.
| Item | Traditional Method | Improved Method | |
| Teacher model inference | Recall | ||
| Rerank | |||
| Total time(Inference Stage) | |||
| Student model training | Extract Features | ||
| Total time(Training Stage) | |||
| Overall | Overall Total |
Appendix F The use of LLMs
We only used GPT for language polishing after completing the first draft, and there were no other scenarios involving the use of Large Language Models (LLMs) in this work.
@@ Appendix F The use of LLMs
| Task | Dataset | Local | Global | ||||
| transition | mining | transition | mining | ||||
| VisualNews | 35.8 | 36.6 | 36.2 | 35.7 | 36.5 | 36.0 | |
| MSCOCO | 83.8 | 80.1 | 82.6 | 77.4 | 74.4 | 73.2 | |
| Fashion200K | 26.1 | 27.6 | 27.9 | 26.1 | 27.6 | 27.9 | |
| WebQA | 92.9 | 95.8 | 96.8 | 92.7 | 95.6 | 96.6 | |
| EDIS | 60.1 | 62.4 | 66.2 | 59.7 | 62.3 | 65.9 | |
| WebQA | 83.4 | 86.1 | 86.7 | 83.0 | 86.0 | 86.3 | |
| VisualNews | 35.2 | 35.0 | 35.7 | 34.8 | 34.9 | 35.6 | |
| MSCOCO | 93.5 | 93.9 | 93.9 | 93.4 | 93.9 | 93.9 | |
| Fashion200K | 27.3 | 28.2 | 26.5 | 27.3 | 28.2 | 26.3 | |
| NIGHTS | 32.4 | 31.9 | 33.3 | 32.3 | 31.9 | 33.2 | |
| OVEN | 51.8 | 55.0 | 57.9 | 46.7 | 47.7 | 53.7 | |
| InfoSeek | 46.5 | 46.2 | 56.8 | 41.4 | 40.9 | 47.7 | |
| FashionIQ | 35.4 | 34.9 | 36.0 | 35.3 | 34.8 | 35.8 | |
| CIRR | 60.3 | 54.2 | 60.5 | 57.7 | 52.8 | 57.2 | |
| OVEN | 71.1 | 75.2 | 73.3 | 65.1 | 67.3 | 67.1 | |
| InfoSeek | 63.7 | 64.0 | 70.1 | 59.3 | 58.9 | 62.7 | |
| Avg. | 56.2 | 56.7 | 58.8 | 54.2 | 54.6 | 56.2 | |
| Methods | Avg. | ||||||||||||||||
| VN | CO | F200 | WQ | ES | WQ | VN | CO | F200 | NS | ON | InS | FQ | CR | ON | InS | ||
| R@5 | R@5 | R@10 | R@5 | R@5 | R@5 | R@5 | R@5 | R@10 | R@5 | R@5 | R@5 | R@10 | R@5 | R@5 | R@5 | ||
| Single model | |||||||||||||||||
| UniIR-BLIP (Wei et al., 2024) | 23 | 75.6 | 25.4 | 79.5 | 50.3 | 79.7 | 21.1 | 88.8 | 27.6 | 33 | 38.7 | 19.7 | 28.5 | 51.4 | 57.8 | 27.7 | 45.5 |
| UniIR-CLIP (Wei et al., 2024) | 42.6 | 77.9 | 17.8 | 84.7 | 59.4 | 78.8 | 42.8 | 92.3 | 17.9 | 32 | 39.2 | 24 | 24.3 | 43.9 | 60.2 | 44.6 | 48.9 |
| MM-Embed (Lin et al., 2024) | 41 | 71.3 | 17.1 | 95.9 | 68.8 | 85 | 41.3 | 90.1 | 18.4 | 32.4 | 42.1 | 42.3 | 25.7 | 50 | 64.1 | 57.7 | 52.7 |
| LamRA-Ret (Liu et al., 2024b) | 41.3 | 75.4 | 28.7 | 85.8 | 62.5 | 81 | 39.3 | 90.4 | 30.4 | 32.1 | 48.4 | 48.7 | 33.1 | 50.5 | 70 | 60 | 54.9 |
[…]
| +Reranker | |||||||||||||||||
| LamRA (Liu et al., 2024b) | 46.9 | 78 | 32.5 | 96.5 | 74.4 | 87.1 | 47.6 | 92.4 | 36.6 | 34.2 | 54 | 58.7 | 37.4 | 59.7 | 72.6 | 74 | 61.4 |
[…]
@@
| UniIR-BLIP (Wei et al., 2024) | 33.0 | 41.0 | 22.4 | 55.8 | 33.0 | 37.0 |
| UniIR-CLIP (Wei et al., 2024) | 32.0 | 45.5 | 27.9 | 67.6 | 48.9 | 44.4 |